OpenStudy (anonymous):

What is the system of linear equation for this solution (1, 2).

OpenStudy (lochana):



OpenStudy (lochana):

so the standard form of a linear equation is \(y = mx + c\)

OpenStudy (lochana):

now you can plug (1,2) in it. and get a relationship with 'm' and 'c'

OpenStudy (anonymous):

b

OpenStudy (anonymous):

y=mx+b not c

OpenStudy (lochana):

\(2 = m + c \ \rightarrow \ c = 2-m \)

OpenStudy (lochana):

ah.. yes. that's right. but it won't be a problem for the final answer.

OpenStudy (lochana):

and the solution will be \( y = mx + 2 -m\)

OpenStudy (lochana):

okay. I agree with @atreyu6s3x6. sorry about that. it is y = mx + b. standard form should be same everywhere.